how do I get y by itself?
x+3y=6

Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

how do I get y by itself?
x+3y=6
---------
Subtract x from both sides to get:
3y = -x+6
Divide thru by 3 to get:
y = (-1/3)x + 2
===========================
Cheers,
Stan H.

Answer by edjones(8007)

x+3y=6
Whatever you do to one side of an equation you must do to the other. This is the essence of algebra.
3y=6-x subtract x from each side.
y=(6-x)/3 divide each side by 3.
.
Ed